What is the variable x after the statement -if2+3=6 if and only if 3+2=5, x = x+1^ - if x = 2 ?

1
Expert's answer
2022-01-24T02:37:30-0500

Solution:

There are two separate statements.

I: 2+3=6 if and only if 3+2=5,

II: x = x+1 if x = 2


Variable x comes in II statement, where x=2 given.